# In-Context Editing: Learning Knowledge from Self-Induced Distributions

### v2

@@ Abstract
In scenarios where language models must incorporate new information efficiently without extensive retraining, traditional fine-tuning methods are prone to overfitting, degraded generalization, and unnatural language generation. To address these limitations, we introduce Consistent In-Context Editing (ICE), a novel approach leveraging the model’s in-context learning capability to optimize towards a contextual distribution rather than a one-hot target. ICE introduces a simple yet effective optimization framework for the model to internalize new knowledge by aligning its output distributions with and without additional context. This method enhances the robustness and effectiveness of gradient-based tuning methods, preventing overfitting and preserving the model’s integrity. We analyze ICE across four critical aspects of knowledge editing: accuracy, locality, generalization, and linguistic quality, demonstrating its advantages. Experimental results confirm the effectiveness of ICE and demonstrate its potential for continual editing, ensuring that the integrity of the model is preserved while updating information.
@@ 1 Introduction
Knowledge editing [53] has emerged as a research area that addresses the challenge of efficiently updating LLM outputs in response to specific queries. It focuses on modifying particular pieces of knowledge in a language model $\mathcal{M}_{\theta}$ using query-response pairs $\{(\mathbf{q}_{i},\mathbf{x}_{i}^{*})\}_{i=1}^{N}$. For instance, given the query ”The president of the US is”, a model trained on outdated data might respond ”Donald Trump”, while the desired up-to-date response would be ”Joe Biden”.
@@
Various strategies have been proposed to address this problem, including constraining the gradient or weights [57; 28] and adopting parameter-efficient fine-tuning approaches [52]. However, these methods still rely on one-hot target distributions, failing to fully mitigate overfitting.
Figure 1: Overview. (a) In-Context Learning: Utilizes context prompts without modifying model parameters, allowing dynamic adaptation but lacking parameter updates. (b) Traditional Fine-Tuning: Minimizes the distance between predictions and a one-hot target ($\delta_{x^{*}}$) using cross-entropy loss ($L_{ft}$), often leading to overfitting. (c) Consistent In-Context Editing (ICE): Adds a contextual loss ($L_{ice}$) to the traditional fine-tuning loss ($L_{ft}$). $L_{ice}$ minimizes the divergence between model outputs with and without a context prompt, aligning the model towards internalizing new knowledge. This helps ICE achieve effective knowledge incorporation while preserving general model stability.
To address these limitations, we introduce Consistent In-Context Editing (ICE), a novel method that learns towards a contextual distribution to effectively internalize new knowledge. Specifically, ICE guides the model’s output distribution $p_{\theta}(\mathbf{x}|\mathbf{q})$ to align with a contextual distribution $p_{\theta}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ induced by a context $\mathbf{c}$ that includes the target knowledge. We minimize the KL divergence between these two distributions, encouraging the model to internalize the new knowledge.
However, the initial contextual distribution may not always perfectly align with the desired update target, so we dynamically adjust it during optimization. We achieve this by combining the contextual loss as a regularization term with the original fine-tuning loss. As the fine-tuning loss is minimized, both the output distribution and the contextual output distribution are guided towards the desired target. The contextual loss, serving as a regularization term, confines the extent of these modifications, thereby ensuring the model’s integrity and preventing unintended degradation. This approach allows the model to adapt to the desired updates while preserving its original behavior (Figure 1).
@@
Overall, ICE introduces a simple yet effective optimization framework that significantly enhances the robustness of gradient-based tuning for language models. At each optimization step, ICE samples in-context sequences and minimizes the difference between outputs with and without context, alongside the fine-tuning loss. This process ensures accurate incorporation of new knowledge, prevents overfitting, and preserves the integrity of previously learned information.
We conduct extensive experiments on four datasets, obtaining promising results across four key dimensions: accuracy, locality, generalization, and linguistic quality. In addition, we evaluate ICE ’s performance without any further design adjustments in continual editing scenarios, where the model undergoes sequential updates, each time with a new knowledge. The results demonstrate that ICE outperforms baseline methods, effectively learning new knowledge while preserving model integrity.
The primary contributions of this paper are: 1) We introduce in-context editing (ICE), a novel knowledge editing approach that learns towards a contextual distribution rather than a one-hot target, offering a more robust alternative to traditional fine-tuning. 2) We develop an optimization framework that refines the target contextual distribution using a gradient-based algorithm, enabling dynamic adaptation of the model to correctly incorporate new knowledge. 3) We provide empirical evidence demonstrating the effectiveness of ICE, showcasing its potential for continual editing by seamlessly integrating new information while preserving the integrity of existing knowledge.
@@ 2.2 Knowledge-Editing Approaches
The first family of methods for knowledge editing keeps the original model frozen while leveraging external tools. Techniques proposed by [31; 32; 27; 34; 56; 47; 46; 22; 4; 38; 45] enhance the model’s adaptability to new knowledge using external memory. Other approaches inject additional parameters into the model to incorporate new knowledge [9; 19; 14; 35; 52]. Additionally, some methods attempt to embed the knowledge directly into prompts to generate post-edit responses, utilizing the model’s in-context learning abilities [55; 6; 24]. However, these methods result in an ever-growing memory/model, which can become problematic over time as knowledge accumulates.
@@
Another line of work, which our method focuses on, involves editing the model’s weights to integrate new knowledge. Approaches include direct direct fine-tuning [57; 23; 33], meta-learning-driven approaches [41; 8; 30; 39], and targeted network editing [10; 7; 25; 5; 36; 50; 42; 20; 13; 15; 16; 1; 11; 51; 49; 26; 12; 44; 37]. These methods aim to incorporate target knowledge and employ various techniques to ensure the locality of the edits. Techniques include constraining the gradient of parameters [57], adopting parameter-efficient approaches [52], and applying statistical constraints on the weights, with notable examples such as ROME [28] and MEMIT [29]. In this paper, we propose tuning the model towards a self-generated distribution instead of a one-hot target.
@@
In-context learning refers to the ability of language models to use information provided in the input context to perform tasks without altering the model’s parameters [2]. Previous research has applied contextual knowledge by prompting the model [55; 6]. To enhance models leveraging in-context learning, various strategies have been explored, such as distilling contextual knowledge using a larger model [40; 18], tuning the model through meta-learning techniques [3], and examining the consistency between context and knowledge [24; 17]. However, these approaches do not modify the model weights. In contrast, our method introduces a novel approach to utilize in-context learning by creating a learning target and framework for model editing, thereby providing an innovative way to integrate contextual information into the model’s knowledge base.
@@ 3.1 Vanilla Fine-Tuning
In our running example, this means we fine-tune the model to assign maximum probability to the sequence ”Joe Biden” given the query ”The president of the US is”. While this approach can effectively update the model’s response for a specific query, it has significant drawbacks. The use of a one-hot target distribution often leads to overfitting, causing the model to degrade, suffer from catastrophic forgetting, or even collapse, resulting in unnatural or repetitive outputs.
@@ 3.2 Fine-Tuning with Sampling
This implies that the model cannot learn and improve on its own without external inputs.
@@ 3.3 In-Context Tuning with Sampling
To address the ineffectiveness of the naive sampling approach in § 3.2, we introduce extra information that guides the model towards a new distribution that aligns with the target, while maintaining similarity to its original distribution. Specifically, we leverage the in-context learning capabilities of language models by prepending context prompts $\mathbf{c}$ to the queries $\mathbf{q}$, where $\mathbf{c}$ includes the new knowledge to be learned. For our example, we can create a context such as ”Joe Biden is the current president of the United States.”, and prepend this context to the query. This induces a new contextual distribution $p_{\theta_{0}}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ that incorporates the desired knowledge through the context, while keeping minimal changes to the model. We can define the loss function as:
@@ 3.4 Consistent In-Context Editing (ICE)
Optimizing $\mathcal{L}_{\text{ICE}}$: The in-context editing loss $\mathcal{L}_{\text{ICE}}$ involves two distributions that depend on the model parameters $\theta$, making direct optimization challenging. To address this, we adopt an iterative, gradient-based approach. At each optimization step $s$, we treat $p_{\theta_{s}}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ as a fixed target distribution (using the current parameters $\theta_{s}$) and update the model parameters to minimize the divergence to $p_{\theta_{s+1}}(\mathbf{x}|\mathbf{q})$. This process is formalized as:
$\mathcal{L}_{\text{ICE}}^{(s)}=D_{\text{KL}}(p_{\theta_{s}}(\mathbf{x}|[\mathbf{c},\mathbf{q}])\;||\;p_{\theta_{s+1}}(\mathbf{x}|\mathbf{q})).$ (7)
@@
Optimizing the Combined Loss $\mathcal{L}$: To optimize the total loss in Equation 6, we sample sequences $\mathbf{x}_{\mathbf{c}}$ from the model conditioned on $[\mathbf{c},\mathbf{q},\mathbf{x}^{*}]$, aiming to maximize the likelihood of the combined sequence $[\mathbf{x}^{*},\mathbf{x}_{\mathbf{c}}]$. This approach helps the model integrate the new knowledge effectively while maintaining fluency and coherence. To prevent the model from drifting too far from the initial parameters (thus preserving unrelated knowledge), we employ gradient clipping techniques inspired by constrained fine-tuning methods [57]. The detailed algorithm is presented in Algorithm 1.
@@ 4.1 Experiment Settings
We evaluate the performance of ICE with four datasets from KnowEdit [53], which are commonly used for knowledge insertion and modification. Detailed statistics on the selected datasets can be seen in Table 1. To ensure a fair comparison, we use Llama2-7b-chat, which is the same model as used in the original survey [53].
@@
We use the metrics from § 2.1 but note one limitation in not penalizing semantically meaningless sentences or repetitive long patterns (§ D.1). Hence we add perplexity as an additional measure, which evaluates a model’s language generation quality. Lower perplexity indicates more natural language. In our case, perplexity can also increase due to the novelty of edited knowledge, so we introduce a normalized perplexity ratio $\text{PPL}_{r}$ to address this (§ D.1). The ratio compares the perplexity of the generation post-target token to that of the combined prompt and target token.
@@ Methods
The contexts $\mathbf{c}$ are given by GPT-4 by summarizing the target knowledge. For layer updates, ROME updates one layer for GPT2 with layer 17 and Llama2 with layer 5. For other baselines, five layers are updated following MEMIT [29], for GPT2 with layers 13,14,15,16,17 and Llama2 with layers 4,5,6,7,8. We follow the usage of other parameters in ROME and MEMIT which have been found to provide the best performance. For FT-M, FT-L, and ICE, the optimization proceeds for a maximum of 25 steps with a learning rate of $47e-$ 4 and 0 weight decay.
@@ 4.2 Main Results
As accuracy increases, the locality tends to decrease due to the inherent perturbations introduced. Furthermore, there tends to be an inverse relationship between model locality and portability; locality implies minimal model changes, whereas portability necessitates the model’s ability to generalize to related knowledge. Despite this trend, ICE not only achieves a near-perfect accuracy comparable to FT-M but also consistently outperforms baseline methods in terms of locality and portability, aligning with the analysis presented in § 3.5. While matching the near perfect accuracy with FT-M, ICE demonstrates consistently better locality and portability than the baseline methods, matching our expectation discussed in § 3.5. Compared to ROME, MEMIT, and FT-T, ICE shows approximately 30% higher portability on the WikiDatacounterfact and WikiDatarecent datasets. This discrepancy highlights that by leveraging in-context learning to adapt to a contextual distribution, ICE achieves better generalization. Additionally, ICE performs over 15% better in terms of locality on both datasets, preserving unrelated knowledge by enhancing the robustness of gradient-based tuning. A minor performance degradation of 99.88% is observed on the WikiBio dataset. This could be attributed to the diversity across datasets, which can introduce slight variations in performance within an acceptable margin.

### v3

@@ Abstract
In scenarios where language models must incorporate new information efficiently without extensive retraining, traditional fine-tuning methods are prone to overfitting, degraded generalization, and unnatural language generation. To address these limitations, we introduce Consistent In-Context Editing (ICE), a novel approach leveraging the model’s in-context learning capability to optimize toward a contextual distribution rather than a one-hot target. ICE introduces a simple yet effective optimization framework for the model to internalize new knowledge by aligning its output distributions with and without additional context. This method enhances the robustness and effectiveness of gradient-based tuning methods, preventing overfitting and preserving the model’s integrity. We analyze ICE across four critical aspects of knowledge editing: accuracy, locality, generalization, and linguistic quality, demonstrating its advantages. Experimental results confirm the effectiveness of ICE and demonstrate its potential for continual editing, ensuring that the integrity of the model is preserved while updating information.
@@ 1 Introduction
Knowledge editing [55] has emerged as a research area that addresses the challenge of efficiently updating LLM outputs in response to specific queries. It focuses on modifying particular pieces of knowledge in a language model $\mathcal{M}_{\theta}$ using query-response pairs $\{(\mathbf{q}_{i},\mathbf{x}_{i}^{*})\}_{i=1}^{N}$. For instance, given the query “The president of the US is”, a model trained on outdated data might respond “Donald Trump”, while the desired up-to-date response would be “Joe Biden”.
@@
Various strategies have been proposed to address this problem, including constraining the gradient or weights [59; 28] and adopting parameter-efficient fine-tuning approaches [54]. However, these methods still rely on one-hot target distributions, failing to fully mitigate overfitting.
Figure 1: Overview. (a) In-Context Learning: Utilizes context prompts without modifying model parameters, allowing dynamic adaptation but lacking parameter updates. (b) Traditional Fine-Tuning: Minimizes the distance between predictions and a one-hot target ($\delta_{x^{*}}$) using cross-entropy loss ($L_{ft}$), often leading to overfitting. (c) Consistent In-Context Editing (ICE): Adds a contextual loss ($L_{ice}$) to the traditional fine-tuning loss ($L_{ft}$). $L_{ice}$ minimizes the divergence between model outputs with and without a context prompt, aligning the model toward internalizing new knowledge. This helps ICE achieve effective knowledge incorporation while preserving general model stability.
To address these limitations, we introduce Consistent In-Context Editing (ICE), a novel method that learns toward a contextual distribution to effectively internalize new knowledge. Specifically, ICE guides the model’s output distribution $p_{\theta}(\mathbf{x}|\mathbf{q})$ to align with a contextual distribution $p_{\theta}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ induced by a context $\mathbf{c}$ that includes the target knowledge. We minimize the KL divergence between these two distributions, encouraging the model to internalize the new knowledge.
However, the initial contextual distribution may not always perfectly align with the desired update target, so we dynamically adjust it during optimization. We achieve this by combining the contextual loss as a regularization term with the original fine-tuning loss. As the fine-tuning loss is minimized, both the output distribution and the contextual output distribution are guided toward the desired target. The contextual loss, serving as a regularization term, confines the extent of these modifications, thereby ensuring the model’s integrity and preventing unintended degradation. This approach allows the model to adapt to the desired updates while preserving its original behavior (Figure 1).
@@
Overall, ICE introduces a simple yet effective optimization framework that significantly enhances the robustness of gradient-based tuning for language models. At each optimization step, ICE samples in-context sequences and minimizes the difference between outputs with and without context, alongside the fine-tuning loss. This process ensures accurate incorporation of new knowledge, prevents overfitting and preserves the integrity of previously learned information.
We conduct extensive experiments on four datasets, obtaining promising results across four key dimensions: accuracy, locality, generalization, and linguistic quality. In addition, we evaluate ICE ’s performance without any adjustments in continual editing scenarios, where the model undergoes sequential updates, each time with a piece of new knowledge. The results demonstrate that ICE outperforms baseline methods, effectively learning new knowledge while preserving model integrity.
The primary contributions of this paper are: 1) We introduce in-context editing (ICE), a novel knowledge editing approach that learns toward a contextual distribution rather than a one-hot target, offering a more robust alternative to traditional fine-tuning. 2) We develop an optimization framework that refines the target contextual distribution using a gradient-based algorithm, enabling dynamic adaptation of the model to correctly incorporate new knowledge. 3) We provide empirical evidence demonstrating the effectiveness of ICE, showcasing its potential for continual editing by seamlessly integrating new information while preserving the integrity of existing knowledge.
@@ 2.2 Knowledge-Editing Approaches
The first family of methods for knowledge editing keeps the original model frozen while leveraging external tools. Techniques proposed by [31; 33; 27; 35; 58; 49; 48; 22; 4; 40; 47] enhance the model’s adaptability to new knowledge using external memory. Other approaches inject additional parameters into the model to incorporate new knowledge [9; 19; 14; 36; 54]. Additionally, some methods attempt to embed the knowledge directly into prompts to generate post-edit responses, utilizing the model’s in-context learning abilities [57; 6; 24]. However, these methods result in an ever-growing memory/model, which can become problematic over time as knowledge accumulates.
@@
Another line of work, which our method focuses on, involves editing the model’s weights to integrate new knowledge. Approaches include direct direct fine-tuning [59; 23; 34], meta-learning-driven approaches [43; 8; 30; 41], and targeted network editing [10; 7; 25; 5; 38; 52; 44; 20; 13; 15; 16; 1; 11; 53; 51; 26; 12; 46; 39; 37]. These methods aim to incorporate target knowledge and employ various techniques to ensure the locality of the edits. Techniques include constraining the gradient of parameters [59], adopting parameter-efficient approaches [54], and applying statistical constraints on the weights, with notable examples such as ROME [28] and MEMIT [29]. In this paper, we propose tuning the model towards a self-generated distribution instead of a one-hot target.
@@
In-context learning refers to the ability of language models to use the information provided in the input context to perform tasks without altering the model’s parameters [2]. Previous research has applied contextual knowledge by prompting the model [57; 6]. To enhance models leveraging in-context learning, various strategies have been explored, such as distilling contextual knowledge [42; 18] and compressing the context into a gist token [32]. However, [42] necessitates maintaining two copies of the model weights, while [32] requires the addition of an extra token to facilitate the injection of new knowledge. There are also methods tuning the model through meta-learning techniques [3], and examining the consistency between context and knowledge [24; 17]. However, these approaches do not modify the model weights. In contrast, our method introduces a novel approach to utilize in-context learning by creating a learning target and framework for model editing, thereby providing an innovative way to integrate contextual information into the model’s knowledge base.
@@ 3.1 Vanilla Fine-Tuning
In our running example, this means we fine-tune the model to assign maximum probability to the sequence “Joe Biden” given the query “The president of the US is”. While this approach can effectively update the model’s response for a specific query, it has significant drawbacks. The use of a one-hot target distribution often leads to overfitting, causing the model to degrade, suffer from catastrophic forgetting, or even collapse, resulting in unnatural or repetitive outputs.
@@ 3.2 Fine-Tuning with Sampling
This implies that the model cannot learn and improve on its own without external inputs, highlighting the necessity for our method, which will be introduced in the following sections.
@@ 3.3 In-Context Tuning with Sampling
To address the ineffectiveness of the naive sampling approach in § 3.2, we introduce extra information that guides the model towards a new distribution that aligns with the target, while maintaining similarity to its original distribution. Specifically, we leverage the in-context learning capabilities of language models by prepending context prompts $\mathbf{c}$ to the queries $\mathbf{q}$, where $\mathbf{c}$ is the new knowledge to be learned. For our example, we can create a context such as “Joe Biden is the current president of the United States.”, and prepend this context to the query. This induces a new contextual distribution $p_{\theta_{0}}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ that incorporates the desired knowledge through the context, while keeping minimal changes to the model. We can define the loss function as:
@@ 3.4 Consistent In-Context Editing (ICE)
Optimizing $\mathcal{L}_{\text{ICE}}$: The in-context editing loss $\mathcal{L}_{\text{ICE}}$ involves two distributions that depend on the model parameters $\theta$, making direct optimization challenging. Directly propagating the loss through both distributions is not desirable, as we aim for a uni-directional optimization: we do not intend to draw $p_{\theta}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ towards $p_{\theta}(\mathbf{x}|\mathbf{q})$. To address this, we adopt an iterative, gradient-based approach. At each optimization step $s$, we treat $p_{\theta_{s}}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ as a fixed target distribution (using the current parameters $\theta_{s}$) and update the model parameters to minimize the divergence to $p_{\theta_{s+1}}(\mathbf{x}|\mathbf{q})$. This process is formalized as:
$\theta_{s+1}^{*}=\mathop{\operator@font arg\min}_{\theta_{s+1}}\mathcal{L}_{\text{ICE}}^{(s)}=\mathop{\operator@font arg\min}_{\theta_{s+1}}D_{\text{KL}}(p_{\theta_{s}}(\mathbf{x}|[\mathbf{c},\mathbf{q}])\;||\;p_{\theta_{s+1}}(\mathbf{x}|\mathbf{q})).$ (7)
@@
Optimizing the Combined Loss $\mathcal{L}$: To optimize the total loss as defined in Equation 6, we sample sequences $\mathbf{x}_{\mathbf{c}}$ from the model conditioned on $[\mathbf{c},\mathbf{q},\mathbf{x}^{*}]$ and maximize the likelihood of the combined sequence $[\mathbf{x}^{*},\mathbf{x}_{\mathbf{c}}]$. This process is equivalent to optimizing the combined loss $\mathcal{L}$. If the sampling is not conditioned on the target, we would be solely optimizing $L_{ICE}$. This approach is algorithmically convenient, and we provide a proof of this equivalence in § A.3. To prevent the model from drifting too far from the initial parameters (thus preserving unrelated knowledge), we employ gradient clipping techniques inspired by constrained fine-tuning methods [59]. The detailed algorithm is presented in Algorithm 1.
@@ 3.5 Discussion
The computational demands of our pipeline can be heavier than vanilla fine-tuning, as it involves multiple sampling steps and depends on GPT-4 for context generation. However, the computational burden may not be as substantial as it appears: 1) Since sampling only necessitates a forward pass of the model, the computational cost is significantly lower than that of training the model. 2) We are considering scenarios with very limited training data, as is the case in the knowledge editing task.
@@ 4.1 Experiment Settings
We evaluate the performance of ICE with four datasets from KnowEdit [55], which are commonly used for knowledge insertion and modification. Detailed statistics on the selected datasets can be seen in Table 1. Among the datasets, the WikiBio dataset does not include related hopping question data necessary for evaluating the portability metric. To ensure a fair comparison, we use Llama2-7b-chat, which is the same model as used in the original survey [55].
@@
We use the metrics from § 2.1 but note one limitation in not penalizing semantically meaningless sentences or repetitive long patterns (§ D.1). Hence we add perplexity as an additional measure, which measures how well the pre-trained model predicts the generated outputs from the fine-tuned models. Assuming the original model is well-trained, the perplexity score reflects the language quality of the fine-tuned model and how far it has drifted. In our case, perplexity can also increase due to the novelty of edited knowledge, so we introduce a normalized perplexity ratio $\text{PPL}_{r}$ to address this (§ D.1). The ratio compares the perplexity of the generation post-target token to that of the combined prompt and target token.
@@ Methods
The contexts $\mathbf{c}$ are given by GPT-4 by summarizing the target knowledge. For layer updates, ROME updates one layer for GPT2 with layer 17 and Llama2 with layer 5. For both ICE and other baselines (FT-M and FT-L), five layers are updated following MEMIT [29], for GPT2 with layers 13,14,15,16,17 and Llama2 with layers 4,5,6,7,8. Results of GPT2 can be found at § D.3. We follow the usage of other parameters in ROME and MEMIT which have been found to provide the best performance. For FT-M, FT-L, and ICE, the optimization proceeds for a maximum of 25 steps with a learning rate of $47e-$ 4 and 0 weight decay. For all results except the ablation study, we used $\lambda=1.0$ for ICE without deliberate tuning.
@@ 4.2 Main Results
As accuracy increases, the locality tends to decrease due to the inherent perturbations introduced. Furthermore, there tends to be an inverse relationship between model locality and portability; locality implies minimal model changes, whereas portability necessitates the model’s ability to generalize to related knowledge. Despite this trend, ICE not only achieves a near-perfect accuracy comparable to FT-M but also consistently outperforms baseline methods in terms of locality and portability, aligning with the analysis presented in § 3.5. While matching the near-perfect accuracy with FT-M, ICE demonstrates consistently better locality and portability than the baseline methods, matching our expectation discussed in § 3.5. Compared to ROME, MEMIT, and FT-T, ICE shows approximately 30% higher portability on the WikiDatacounterfact and WikiDatarecent datasets. This discrepancy highlights that by leveraging in-context learning to adapt to a contextual distribution, ICE achieves better generalization. Additionally, ICE performs over 15% better in terms of locality on both datasets, preserving unrelated knowledge by enhancing the robustness of gradient-based tuning. A minor performance degradation of 99.88% is observed on the WikiBio dataset. This could be attributed to the diversity across datasets, which can introduce slight variations in performance within an acceptable margin.
@@
Table 5: Ablation results for different values of $\lambda$.
ZsRE $\textbf{WikiData}_{recent}$
[…]
0.6 99.71 50.65 59.54 584.84 561.29 99.93 58.28 46.93 589.77 179.4
0.8 99.81 51.59 58.82 582.91 281.14 99.95 59.12 47.36 591.84 142.03
1.0 100.00 55.52 56.97 562.70 15.50 100.00 61.02 46.39 585.58 34.08
1.2 99.87 52.08 58.54 581.04 324.47 99.98 59.69 47.51 589.49 280.32
1.4 99.90 51.91 58.18 584.17 287.25 100.00 59.96 46.53 591.98 154.77